A merchant wishes to mix candy worth $5 per lb with 40lb of candy worth $2 per pound to get a mixture that can be sold for $3 per lb. How many pounds of $5 candy should be used?
Answer by checkley75(3666)

5X+2*40=3(40+X)
5X+80=120+3X
5X-3X=120-80
2X=40
X=40/2
X=20 POUNDS OF $5 CANDY IS NEEDED.
PROOF
5*20+2*40=3*60
100+80=180
180=180